The market price of a stock is $22.32 and it just paid a dividend of $1.88. The required rate of return is 11.16%. What is the expected growth rate of the dividend?
Answer #1

D0 = $1.88

D1 = D0(1+g)

= 1.88(1+g)

P0 = D1 / (k - g)

22.32 = 1.88(1+g) / (0.1116 - g)

22.32(0.1116 - g) = 1.88(1+g)

2.490912 - 22.32g = 1.88 + 1.88g

2.490912 - 1.88 = 1.88g + 22.32g

0.610912 = 24.2g

g = 0.610912 / 24.2

g = 0.02524 or 2.524%

Therefore,

Expected Growth Rate = 2.524%